11. The battery life of a certain battery is normally distributed with a mean of 90 days and a standard deviation of 3 days.

For each of the following questions, construct a normal distribution curve and provide the answer.

a)    About what percent of the products last between 87 and 93 days?

b)    About what percent of the products last 84 or less days?

For each of the following questions, use the standard normal table and provide the answer.

c) About what percent of the products last between 89 and 94 days?

d) About what percent of the products last 95 or more days? 




1
Expert's answer
2022-04-13T16:05:03-0400

a) The range is one standard deviation below and above the average value. So according to the 68–95–99.7 rule, the answer is 68%

b) These products lie under two standard deviations, so their percentage according to the 68–95–99.7 rule is (100 - 95)/2 = 2.5%

c) Let's convert 89 and 94 days to standatd normal distribution: z1 = (89-90)/3 = -0.33, 94 and z2 = (94-90)/3 = 1.33, therefore

"p = \Phi(1.33) - \Phi(-0.33) = 0.90824 - 0.37070 = 0.53754" or 53.8%

d) 95 days when converted to standard normal distribution corresponds to z = (95-90)/3 = 1.66, therefore "p = 1-\Phi(1.66) = 1 - 0.95154=0.04846" or 4.8%